# The ECME Algorithm Using Factor Analysis for DOA Estimation in Nonuniform Noise

### v2

@@ 1 Introduction
Factor analysis (FA) has been a tool in psychology for over one hundred years [1] and its purpose is to extract or estimate the low-rank covariance matrix from a measurement covariance matrix. To this end, several techniques have been developed, but the only one available, which provides efficient and consistent estimators, is maximum likelihood FA (MLFA) [2, 3, 4]. Unfortunately, these estimators cannot be expressed in closed form, which leads to a variety of iterative approaches [5]. In particular, one of the most promising is the Expectation-Maximization (EM) algorithm [6].
@@
The EM algorithm regards observations or measurements as the incomplete data and designs the complete data containing missing (latent or unobserved) data. The complete data is so manipulable that its log-likelihood function (LLF) is more tractable than the incomplete-data or actual LLF. Hence, the EM algorithm uses the complete-data LLF to efficiently update the estimates of parameters, which can still increase the actual likelihood. To be specific, each iteration of the EM algorithm consists of an expectation step (E-step) and a maximization step (M-step). However, the M-step requires maximizing the expected complete-data LLF provided by the E-step, which causes two shortcomings: 1) slow convergence, and 2) difficult maximization in numerous complex cases. To make up for both shortcomings, the Expectation/Conditional Maximization Either (ECME) algorithm, an extension of the EM algorithm, has been proposed [7] and later applied to MLFA [8]. However, a conditional M-step (CM-step) at its each iteration utilizes Newton-Raphson, instead of an explicit formula, to update the estimates of partial parameters by the actual LLF and thus is complex as well as unstable.
@@
In addition to psychometrics, econometrics, statistics and so on, FA recently finds applications in array signal processing, especially including direction of arrival (DOA) estimation in unknown nonuniform noise [9, 10]. The nonuniform noise assumption can let the signal received at an array be modeled by FA so that MLFA is used to estimate the signal subspace for the subspace based DOA estimation [11]. For this purpose, an iterative approach, called Iterative ML Subspace Estimation (IMLSE), is proposed in [9]. This method sequentially updates the estimates of parameters at each iteration by explicit formulas and is computationally attractive. However, Stoica and Babu have shown that the IMLSE method cannot guarantee its monotonicity and may fail to converge [10]. Afterwards, they propose a novel iterative approach called Factor Analysis for Anisotropic Noise (FAAN), which has excellent convergence properties.
@@
In this article, the ECME algorithm is designed again for MLFA, which has almost the same computational complexity at each iteration as the FAAN method. However, numerical results show that the ECME algorithm yields faster stable convergence and is computationally more efficient. Importantly, the signal subspace estimated by the ECME algorithm can be employed for the subspace based DOA estimation in unknown nonuniform noise.
@@
Notations: $(\cdot)^{T}$ and $(\cdot)^{H}$ mean the transpose and conjugate transpose operators, respectively. $\mathrm{Det}(\boldsymbol{\Sigma})$, $\mathrm{Tr}\{\boldsymbol{\Sigma}\}$, $\boldsymbol{\Sigma}^{1/2}$, and $\boldsymbol{\Sigma}^{-1}$ stands for the determinant, trace, square root, and inverse of square matrix $\boldsymbol{\Sigma}$. $\Sigma_{i,j}$ is the element in the $i$th row and $j$th column of $\boldsymbol{\Sigma}$. $\boldsymbol{\Sigma}\succ\mathbf{0}_{N}$ and $\boldsymbol{\Sigma}\succeq\mathbf{0}_{N}$ denote that the $N\times N$ matrix $\boldsymbol{\Sigma}$ is positive definite and positive semi-definite, respectively. $\mathbf{I}_{N}$ is the identity matrix of order $N$ and $\jmath$ is the imaginary unit. $\mathcal{C}(\boldsymbol{\Sigma})$ denotes the column space of $\boldsymbol{\Sigma}$. $(\cdot)^{(k)}$ is an iterative value obtained at the $k$th iteration.
@@ 2 Problem Description
We consider a uniform linear array (ULA) consisting of $N$ isotropic sensors for analytical simplicity. $M<N$ narrow-band source signals are assumed to impinge upon the array from far field and the distance between arbitrary two adjacent sensors is the half-wavelength of the source signals. Thus, the composite signal received at this ULA can be expressed as
@@
where $\theta_{m}\in(0,\pi)$ represents the DOA of the $m$th source signal, $\mathbf{a}(\theta_{m})=[1~\cdots~e^{-(N-1)\jmath\pi\cos(\theta_{m})}]^{T}\in\mathbb{C}^{N}$ is the steering vector of the $m$th source, and $\mathbf{A}=[\mathbf{a}(\theta_{1})~\cdots~\mathbf{a}(\theta_{M})]$ is the array manifold matrix. Additionally, $\mathbf{s}(t)=[s_{1}(t)~\cdots~s_{M}(t)]^{T}\in\mathbb{C}^{M}$ is the source signal vector, $\mathbf{v}(t)$ is a spatially nonuniform Gaussian white noise vector with mean zero and diagonal covariance matrix $\mathbf{Q}=\mathrm{diag}\{\sigma^{2}_{1},\dots,\sigma^{2}_{N}\}\succ\mathbf{0}_{N}$, i.e., $\mathbf{v}(t)\sim\mathcal{CN}(\mathbf{0},\mathbf{Q})$.
When the stochastic source signal model is adopted, $\mathbf{s}(t)\sim\mathcal{CN}(\mathbf{0},\mathbf{P})$, where $\mathbf{P}\succeq\mathbf{0}_{M}$ is the source covariance matrix. In (1), $\mathbf{s}(t)$ and $\mathbf{v}(t)$ are assumed to be mutually uncorrelated, which leads to $\mathbf{y}(t)\sim\mathcal{CN}(\mathbf{0},\mathbf{C})$ with $\mathbf{C}=\mathbf{A}\mathbf{P}\mathbf{A}^{H}+\mathbf{Q}$ [11]. In practice, we first sample $\mathbf{y}(t)$ and obtain independent and identically distributed snapshots as follows:
@@
From the perspective of ML direction finding [12], we need to write the LLF of $\mathbf{Y}$ with $\mathbf{Y}=[\mathbf{y}(1)~\cdots~\mathbf{y}(L)]$ as
$l(\boldsymbol{\Phi})$ $=$ $\ln p(\mathbf{Y};\boldsymbol{\Phi})={\sum}_{t=1}^{L}\ln p\big{(}\mathbf{y}(t);\boldsymbol{\Phi}\big{)}$
$=$ $-LN\ln(\pi)-L\Big{(}\ln\mathrm{Det}(\mathbf{C})+\mathrm{Tr}\{\hat{\mathbf{R}}\mathbf{C}^{-1}\}\Big{)}$
where $\boldsymbol{\Phi}=(\boldsymbol{\theta},\mathbf{P},\mathbf{Q})$ with $\boldsymbol{\theta}=(\theta_{1},\dots,\theta_{M})$, $p(\cdot)$ is the probability density function, and $\hat{\mathbf{R}}=(1/L)\sum_{t=1}^{L}\mathbf{y}(t)\mathbf{y}^{H}(t)\succ\mathbf{0}_{N}$ [^1] In this article, we assume $\hat{\mathbf{R}}\succ\mathbf{0}_{N}$, which requires $L\geq N$.. Importantly, we need to solve the equivalent ML direction finding problem [13, 14]:
@@
In order to solve problem (3), the ECME algorithm has been designed in the previous work [15]. In next sections, we will design the ECME algorithm using FA for the subspace based DOA estimation.
@@ 3.1 MLFA Estimation Problem
It is well-known that the subspace based DOA estimation only requires the signal subspace $\mathcal{C}(\mathbf{A})$ [11]. As a result, let $\mathbf{C}$ be the population covariance matrix in FA [3]
@@
where all elements in $\mathbf{S}\in\mathbb{C}^{N\times M}$ are free parameters (factor loadings) and hence $\mathbf{SS}^{H}$ must be able to equal any $\mathbf{APA}^{H}$. Here, we also use $\mathbf{S}$ to estimate $\mathcal{C}(\mathbf{A})$. Eq. (4) makes problem (3) become the classic MLFA estimation problem [2, 4, 3, 5]:
@@
For solving problem (5), the IMLSE method utilizes a fixed-point iteration and sequentially (not simultaneously) updates the estimates of $\mathbf{S}$ and $\mathbf{Q}$ at each iteration by two explicit formulas [9]. However, Stoica and Babu have shown that the IMLSE method cannot guarantee its monotonicity and may fail to converge [10]. Hence, they propose the FAAN method with excellent convergence properties. We outline the FAAN method in the next subsection.
@@ 3.2 FAAN Method
When a direct minimization over all free parameters is very difficult, an iterative (coordinate descent) approach or alternating minimization can always be adopted [16]. As mentioned before, the FAAN method has been proposed for problem (5) and its each iteration consists of two steps. Specifically, since $\mathbf{C}=\mathbf{SS}^{H}+\mathbf{Q}$ makes updating the estimates of $\mathbf{S}$ and $\mathbf{Q}$ difficult, a new parametrization of $\mathbf{C}$ is designed as
@@
where $\boldsymbol{\Lambda}\succeq\mathbf{0}_{M}$, $\mathbf{U}\boldsymbol{\Lambda}\mathbf{U}^{H}$ is the eigen-decomposition of $\mathbf{Q}^{-1/2}\mathbf{SS}^{H}\mathbf{Q}^{-1/2}$, and $\mathbf{S}=\mathbf{Q}^{1/2}\mathbf{U}\boldsymbol{\Lambda}^{1/2}$. Thus, problem (5) is changed to
@@
According to (3.2), the first step at the $k$th iteration of the FAAN method simultaneously obtains $\mathbf{U}^{(k)}$ and $\boldsymbol{\Lambda}^{(k)}$ by
@@
where $\tilde{\mathbf{R}}^{(k)}=[\mathbf{Q}^{(k-1)}]^{-1/2}\hat{\mathbf{R}}[\mathbf{Q}^{(k-1)}]^{-1/2}$. This leads to
@@
where the $\lambda_{m}^{(k)}$’s are the eigenvalues of $\tilde{\mathbf{R}}^{(k)}$ and $\lambda_{1}^{(k)}\geq\cdots\geq\lambda_{N}^{(k)}$, $\mathbf{B}^{(k)}$ is composed of the $M$ principal eigenvectors. Moreover, the operator $(\cdot)_{+}$ replaces the negative value with zero and leaves the nonnegative value unchanged.
The second step at the $k$th iteration of the FAAN method sequentially (not simultaneously) obtains $\sigma_{1}^{(k)},\dots,\sigma_{N}^{(k)}$ by
@@
where $\boldsymbol{\Gamma}^{(k)}=\big{(}\mathbf{U}^{(k)}\boldsymbol{\Lambda}^{(k)}[\mathbf{U}^{(k)}]^{H}+\mathbf{I}_{N}\big{)}^{-1}$ [^2] We do not adopt $\boldsymbol{\Gamma}^{(k)}=\mathbf{I}_{N}+\mathbf{U}^{(k)}\big{(}[\boldsymbol{\Lambda}^{(k)}]^{-1}-\mathbf{I}_{M}\big{)}[\mathbf{U}^{(k)}]^{H}$ since numerical results show its low accuracy.. This leads to
$\sigma_{n}$ $=$ $\frac{1}{2}\Big{(}b_{n}^{(k)}+\sqrt{[b_{n}^{(k)}]^{2}+4c_{n}^{(k)}}\Big{)}>0,n=1,\dots,N$
where $\Re(\cdot)$ denotes the real part of the complex value, $b_{n}^{(k)}=\sum_{i\neq n}\Re\{\hat{R}_{i,n}\Gamma_{i,n}^{(k)}\}/\sigma_{i}$ and $c_{n}^{(k)}=\hat{R}_{n,n}\Gamma_{n,n}^{(k)}>0$. The details of the FAAN method are in Algorithm 1.
@@
4:  Obtain $\mathbf{Q}^{(k)}$ using (3.2) $K_{2}$ times.
@@ 4 ECME Algorithm
The ECME algorithm has been designed for problem (5) in [8] and yields rapider convergence than the EM algorithm in [17]. However, a CM-step at its each iteration utilizes Newton-Raphson, instead of an explicit formula, to update the estimate of $\mathbf{Q}$ by the actual or negative LLF $f(\mathbf{S},\mathbf{Q})=\ln\mathrm{Det}(\mathbf{C})+\mathrm{Tr}\{\hat{\mathbf{R}}\mathbf{C}^{-1}\}$ in (5), leading to that this step is complex and the monotonicity cannot be guaranteed. In this section, we again design the ECME algorithm with excellent convergence.

### v3

@@ 1 Introduction
Factor analysis (FA) has been a tool in psychology for over one hundred years [spearman] and its purpose is to extract or estimate the low-rank covariance matrix from a measurement covariance matrix. To this end, several techniques have been developed, but the only one available, which provides an efficient and consistent estimator, is maximum likelihood FA (MLFA) [lawley, basilevsky, mulaik]. Unfortunately, this ML based estimator cannot be expressed in closed form, which leads to a variety of iterative approaches [jöreskog]. In particular, one of the most promising is the Expectation-Maximization (EM) algorithm [dempster, rubin].
@@
The EM algorithm regards observations or measurements as the incomplete data and introduces the complete data containing missing (latent or unobserved) data. The complete data is so manipulable that its log-likelihood function (LLF) is more tractable than the incomplete-data or actual LLF. Accordingly, the EM algorithm only uses the complete-data LLF to efficiently update the estimates of parameters at each iteration, which can still increase the actual likelihood. To be specific, each iteration of the EM algorithm consists of an expectation step (E-step) and a maximization step (M-step). However, the M-step requires maximizing the expected complete-data LLF provided by the E-step, which causes two shortcomings: 1) slow convergence, and 2) difficult maximization in numerous complex cases. To make up for both shortcomings, the Expectation/Conditional Maximization Either (ECME) algorithm, an extension of the EM algorithm, has been proposed [liu] and later applied to MLFA [liu2]. However, a conditional M-step (CM-step) at its each iteration utilizes Newton-Raphson, instead of an explicit formula, to update the estimates of partial parameters by the actual LLF and thus is complex as well as unstable.
@@
In addition to psychometrics, econometrics, statistics and so on, FA recently finds applications in array signal processing, especially including direction of arrival (DOA) estimation in unknown nonuniform noise [liao, stoica]. The nonuniform noise assumption can let the signal received at an array be modeled by FA so that MLFA is employed to estimate the signal subspace for the subspace based DOA estimation [krim]. For this purpose, the iterative approach called Iterative ML Subspace Estimation (IMLSE) is proposed in [liao]. This method sequentially updates the estimates of parameters at each iteration by two explicit formulas and is computationally attractive. However, Stoica and Babu have shown that the IMLSE method cannot guarantee its monotonicity and may fail to converge [stoica]. Afterwards, they propose the novel iterative approach called Factor Analysis for Anisotropic Noise (FAAN), which has excellent convergence properties.
@@
In this article, the ECME algorithm is designed again for MLFA by introducing new complete data, which can thus use two explicit formulas to sequentially update the estimates of parameters at each iteration and have excellent convergence properties. Theoretical analysis shows that the ECME algorithm has almost the same computational complexity at each iteration as the FAAN method. However, numerical results show that the ECME algorithm yields faster stable convergence and the convergence to the global optimum is easier. Importantly, MLFA is not the best choice for the subspace based DOA estimation in unknown nonuniform noise.
@@
Notations: $(\cdot)^{T}$ and $(\cdot)^{H}$ mean the transpose and conjugate transpose operators, respectively. $\mathrm{Det}(\boldsymbol{\Sigma})$, $\mathrm{Tr}\{\boldsymbol{\Sigma}\}$, $\boldsymbol{\Sigma}^{1/2}$, and $\boldsymbol{\Sigma}^{-1}$ stand for the determinant, trace, square root, and inverse of square matrix $\boldsymbol{\Sigma}$. $\Sigma_{i,j}$ is the element in the $i$th row and $j$th column of $\boldsymbol{\Sigma}$. $\boldsymbol{\Sigma}\succ\mathbf{0}_{N}$ and $\boldsymbol{\Sigma}\succeq\mathbf{0}_{N}$ denote that the $N\times N$ matrix $\boldsymbol{\Sigma}$ is positive definite and positive semi-definite, respectively. $\mathbf{I}_{N}$ is the identity matrix of order $N$ and $\jmath$ is the imaginary unit. $\mathcal{C}(\boldsymbol{\Sigma})$ denotes the column space of $\boldsymbol{\Sigma}$. $\theta^{(k)}$ is the iterative value of $\theta$ obtained at the $k$th iteration of an algorithm.
@@ 2 Problem Description
We consider a uniform linear array (ULA) consisting of $N$ isotropic sensors for analytical simplicity. M(M<N)M~(M<N) narrow-band source signals are assumed to impinge upon the array from far field and the distance between arbitrary two adjacent sensors is the half-wavelength of the source signals. Thus, the composite signal received at this ULA can be expressed as
@@
where $\theta_{m}\in(0,\pi)~(\mathrm{radian})$ represents the DOA of the $m$th source signal, $\mathbf{a}(\theta_{m})=[1~e^{-\jmath\pi\cos(\theta_{m})}~\cdots~e^{-(N-1)\jmath\pi\cos(\theta_{m})}]^{T}\in\mathbb{C}^{N}$ is the steering vector of the $m$th source signal, and $\mathbf{A}=[\mathbf{a}(\theta_{1})~\cdots~\mathbf{a}(\theta_{M})]$ is the array manifold matrix. For convenience, let $\theta_{1}<\cdots<\theta_{M}$. Additionally, $\mathbf{s}(t)=[s_{1}(t)~\cdots~s_{M}(t)]^{T}\in\mathbb{C}^{M}$ is the source signal vector, $\mathbf{v}(t)$ is the spatially nonuniform Gaussian white noise vector with mean zero and diagonal covariance matrix $\mathbf{Q}=\mathrm{diag}\{\sigma^{2}_{1},\dots,\sigma^{2}_{N}\}\succ\mathbf{0}_{N}$, i.e., $\mathbf{v}(t)\sim\mathcal{CN}(\mathbf{0},\mathbf{Q})$.
When the stochastic source signal model is adopted, $\mathbf{s}(t)\sim\mathcal{CN}(\mathbf{0},\mathbf{P})$, where $\mathbf{P}\succeq\mathbf{0}_{M}$ is the source covariance matrix. In (1), $\mathbf{s}(t)$ and $\mathbf{v}(t)$ are assumed to be mutually uncorrelated, which leads to $\mathbf{y}(t)\sim\mathcal{CN}(\mathbf{0},\mathbf{C})$ with $\mathbf{C}=\mathbf{A}\mathbf{P}\mathbf{A}^{H}+\mathbf{Q}\succ\mathbf{0}_{N}$ [krim]. In practice, we first sample $\mathbf{y}(t)$ and obtain independent and identically distributed snapshots as follows:
@@
From the processing perspective of ML direction finding [kay], we need to write the LLF of $\mathbf{Y}$ with $\mathbf{Y}=[\mathbf{y}(1)~\cdots~\mathbf{y}(L)]$ as
$l(\boldsymbol{\Phi})$ $=$ $\ln p(\mathbf{Y};\boldsymbol{\Phi})={\sum}_{t=1}^{L}\ln p\big(\mathbf{y}(t);\boldsymbol{\Phi}\big)$ (3)
$=$ $a-L\Big(\ln\mathrm{Det}(\mathbf{C})+\mathrm{Tr}\{\hat{\mathbf{R}}\mathbf{C}^{-1}\}\Big)$
where $a$ is a constant, $\boldsymbol{\Phi}=(\boldsymbol{\theta},\mathbf{P},\mathbf{Q})$ with $\boldsymbol{\theta}=(\theta_{1},\dots,\theta_{M})$, $p(\cdot)$ is the probability density function (PDF), and $\hat{\mathbf{R}}=(1/L)\sum_{t=1}^{L}\mathbf{y}(t)\mathbf{y}^{H}(t)\succ\mathbf{0}_{N}$ [^1] In this article, we assume $\hat{\mathbf{R}}\succ\mathbf{0}_{N}$, which requires $L\geq N$. is the snapshot covariance matrix. Importantly, we need to solve the equivalent ML direction finding problem [jaffer, stoica2, pesavento]:
@@ 3.1 MLFA Estimation Problem
It is well-known that the subspace based DOA estimation only requires the signal subspace $\mathcal{C}(\mathbf{A})$ [krim]. As a result, let $\mathbf{C}$ be the population covariance matrix in FA [basilevsky]
@@
where all the $NM$ elements in $\mathbf{S}\in\mathbb{C}^{N\times M}$ are free parameters (factor loadings) and hence $\mathbf{SS}^{H}$ must be able to equal any $\mathbf{APA}^{H}$. In this article, we also use $\mathbf{S}$ to estimate $\mathcal{C}(\mathbf{A})$. Eq. (5) makes problem (4) become the classic MLFA estimation problem [lawley, basilevsky, mulaik, jöreskog]:
@@
For solving problem (6), the IMLSE method utilizes a fixed-point iteration and sequentially (not simultaneously) updates the estimates of $\mathbf{S}$ and $\mathbf{Q}$ at each iteration by two explicit formulas [liao]. However, Stoica and Babu have shown that the IMLSE method cannot guarantee its monotonicity and may fail to converge [stoica]. Hence, they propose the FAAN method with excellent convergence properties. We outline the FAAN method in the next subsection.
@@ 3.2 FAAN Method
When a direct minimization over all free parameters is very difficult, an iterative (coordinate descent) approach or alternating minimization can always be adopted [bezdek]. As mentioned before, the FAAN method has been proposed for problem (6) and its each iteration consists of two steps. Specifically, since $\mathbf{C}=\mathbf{SS}^{H}+\mathbf{Q}$ makes updating the estimates of $\mathbf{S}$ and $\mathbf{Q}$ difficult, a new parametrization of $\mathbf{C}$ is designed as
@@
where $\boldsymbol{\Lambda}\succeq\mathbf{0}_{M}$, $\mathbf{U}\boldsymbol{\Lambda}\mathbf{U}^{H}$ is the eigen-decomposition of $\mathbf{Q}^{-1/2}\mathbf{SS}^{H}\mathbf{Q}^{-1/2}$, and $\mathbf{S}=\mathbf{Q}^{1/2}\mathbf{U}\boldsymbol{\Lambda}^{1/2}$ for convenience. Thus, problem (6) is changed to
@@
According to (3.2), given $\mathbf{Q}=\mathbf{Q}^{(k-1)}\succ\mathbf{0}_{N}$, the first step at the $k$th iteration of the FAAN method simultaneously obtains $\mathbf{U}^{(k)}$ and $\boldsymbol{\Lambda}^{(k)}$ by
@@
where $\tilde{\mathbf{R}}^{(k)}=[\mathbf{Q}^{(k-1)}]^{-1/2}\hat{\mathbf{R}}[\mathbf{Q}^{(k-1)}]^{-1/2}\succ\mathbf{0}_{N}$. This leads to
@@
where the $\lambda_{m}^{(k)}$’s are the eigenvalues of $\tilde{\mathbf{R}}^{(k)}$ and $\lambda_{1}^{(k)}\geq\cdots\geq\lambda_{N}^{(k)}>0$, the $m$th column of $\mathbf{B}^{(k)}$ is the eigenvector corresponding to $\lambda_{m}^{(k)}$. Moreover, the operator $(b)_{+}$ replaces the negative value of real number $b$ with zero and leaves the nonnegative value unchanged.
The second step at the $k$th iteration of the FAAN method sequentially (one by one) obtains $\sigma_{1}^{(k)},\dots,\sigma_{N}^{(k)}$ by the $N$ problems:
@@
where $n=1,\dots,N$ and $\boldsymbol{\Gamma}^{(k)}=\big(\mathbf{U}^{(k)}\boldsymbol{\Lambda}^{(k)}[\mathbf{U}^{(k)}]^{H}+\mathbf{I}_{N}\big)^{-1}\succ\mathbf{0}_{N}$ [^2] We do not adopt $\boldsymbol{\Gamma}^{(k)}=\mathbf{I}_{N}+\mathbf{U}^{(k)}\big([\boldsymbol{\Lambda}^{(k)}]^{-1}-\mathbf{I}_{M}\big)[\mathbf{U}^{(k)}]^{H}$ since numerical results show its low accuracy.. This leads to
$\sigma_{n}=\frac{1}{2}\Big(b_{n}^{(k)}+\sqrt{[b_{n}^{(k)}]^{2}+4c_{n}^{(k)}}\Big)>0$ (12)
where $n=1,\dots,N$, $\Re(b)$ denotes the real part of complex number $b$, $b_{n}^{(k)}=\sum_{i\neq n}\Re\{\hat{R}_{i,n}\Gamma_{i,n}^{(k)}\}/\sigma_{i}$ and $c_{n}^{(k)}=\hat{R}_{n,n}\Gamma_{n,n}^{(k)}>0$. The details of the FAAN method are in Algorithm 1.
@@
4:  Obtain $\mathbf{Q}^{(k)}$ using (12) $K_{2}$ times.
@@ 4 ECME Algorithm
The ECME algorithm has been designed for problem (6) in [liu2] and yields rapider convergence than the EM algorithm designed in [rubin]. However, a CM-step at its each iteration utilizes Newton-Raphson, instead of an explicit formula, to update the estimate of $\mathbf{Q}$ by the objective function or negative LLF $f(\mathbf{S},\mathbf{Q})=\ln\mathrm{Det}(\mathbf{C})+\mathrm{Tr}\{\hat{\mathbf{R}}\mathbf{C}^{-1}\}$ in problem (6), leading to that this step is complex and the monotonicity cannot be guaranteed. In this section, we again design the ECME algorithm by introducing new complete data, which can thus use two explicit formulas to sequentially update the estimates of $\mathbf{S}$ and $\mathbf{Q}$ at each iteration and have excellent convergence properties.